OpenStudy (anonymous):

Divide.Write your answer as a fraction or a mixed number in simplest form. 3 1/2 divided by 2 2/3. i got 9/8

OpenStudy (anonymous):

3 (1/2) = 3/2 if like this u are right 3 1/2 = 7/2 the answer is 7/2 x 3/8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

there are 13 books on a shelf. 8 of these books are new. (a.) what is the ratio of used books to new books? (b.) what is the ratio of used books to all books?